Point D(6,-3)is translated 3 units to the left and 1 units down to form the point D' . What are the resulting coordinates of D' ?
1
Expert's answer
2012-09-27T11:48:30-0400
D'(x) = D(x) - 3 = 6 - 3 = 3;
D'(y) = D(y) - 1 = -3 - 1 = -4.

So, D' has coordinates of D'(3,-4).